OpenStudy (anonymous):

what percent of 200 is 161.7

OpenStudy (anonymous):

161.7/2 %

OpenStudy (anonymous):

161.7/200=.8085x100=80.85

OpenStudy (anonymous):

\[\frac{161.7}{200}\times 100=80.85\]

OpenStudy (anonymous):

thanks guys